# Improving SDSS Cosmological Constraints through $\beta$-Skeleton Weighted Correlation Functions

### v1

@@ V Concluding Remarks
We have measured three MCFs, each weighted by: i) the number of neighboring galaxies around each galaxy, $N_{\rm con}$; ii) the average distance of each galaxy from its surrounding neighbors, $\bar{D}_{\rm nei}$; iii) the reciprocal of the average distance of each galaxy from its surrounding neighbors, $1/\bar{D}_{\rm nei}$. Through a comparison of measurements and the calculation of corresponding $\chi^{2}$ statistics, we observe a substantial improvement in the constraints on the cosmological parameter $\Omega_{m}$ by conducting a joint analysis of the standard 2PCF and all three MCFs for different weights.
@@ Acknowledgments
This work is supported by National SKA Program of China (2020SKA0110401, 2020SKA0110402, 2020SKA0110100), the National Key R&D Program of China (2020YFC2201600, 2018YFA0404504, 2018YFA0404601), the National Science Foundation of China (11890691, 12203107, 12073088, 12373005), the China Manned Space Project with No. CMS-CSST-2021 (A02, A03, B01), the Guangdong Basic and Applied Basic Research Foundation (2019A1515111098), and the 111 project of the Ministry of Education No. B20019. We also wish to acknowledge the Beijing Super Cloud Center (BSCC) and Beijing Beilong Super Cloud Computing Co., Ltd (http://www.blsc.cn/) for providing HPC resources that have significantly contributed to the research results presented in this paper.

### v3

@@ V Concluding Remarks
We have measured three MCFs, each weighted by i) the number of neighboring galaxies around each galaxy, $N_{\rm con}$; ii) the average distance of each galaxy from its surrounding neighbors, $\bar{D}_{\rm nei}$; and iii) the reciprocal of the average distance of each galaxy from its surrounding neighbors, $1/\bar{D}_{\rm nei}$. Through a comparison of measurements and the calculation of corresponding $\chi^{2}$ statistics, we observe a substantial improvement in the constraints on the cosmological parameter $\Omega_{m}$ by conducting a joint analysis of the standard 2PCF and all three MCFs for different weights.
@@ Acknowledgments
This work is supported by National SKA Program of China (Grants No. 2020SKA0110401, No. 2020SKA0110402, No. 2020SKA0110100), the National Key R and D Program of China (Grants No. 2020YFC2201600, No. 2018YFA0404504, No. 2018YFA0404601), the National Science Foundation of China (Grants No. 11890691, No. 12203107, 12073088, No. 12373005), the China Manned Space Project with No. CMS-CSST-2021 (( Projects No. A02, No. A03, No. B01), the Guangdong Basic and Applied Basic Research Foundation (2019A1515111098), and the 111 project of the Ministry of Education No. B20019. We also wish to acknowledge the Beijing Super Cloud Center (BSCC) and Beijing Beilong Super Cloud Computing Co., Ltd (http://www.blsc.cn/) for providing HPC resources that have significantly contributed to the research results presented in this paper.
APPENDIX
To ensure that the accuracy of COLA simulation is sufficient for this analysis, here we have made a comparison between the statistics measured in COLA and GADGET Springel et al. (2001b); Springel (2005); Springel et al. (2021) simulations. GADGET is a massively parallel TreeSPH code for accurate N-body/SPH simulations Xu (1995), which can simulate a collisionless fluid by the N-body method or an ideal gas with smoothed particle hydrodynamics (SPH). (Ding et al., 2024) combines the subhalo abundance matching (SHAM) procedure with COLA and GADGET simulations to create mocks for BOSS CMASS NGC galaxies by fitting the correlation function at $s\in[4,20]~{}h^{-1}\rm Mpc$ within the redshift interval $z\in[0.45,0.55]$. Below we compare the clustering properties of the COLA, GADGET and PATCHY mocks with the observed data.
To ensure the reliability of our analysis, the COLA and GADGET mocks used the same cosmological parameters $\{\Omega_{m}=0.2951,\Omega_{\Lambda}=0.7049,\Omega_{b}=0.0468,w=-1.0,\sigma_{8}=0.80,n_{s}=0.96\}$ for simulating DM particles and ran simulations under identical conditions with $1024^{3}$ particles in a box measuring $800~{}h^{-1}\rm Mpc$.
[…]
Figure 12: Similar to Fig. 11, but for the normalized MCFs, $s^{2}\hat{W}_{0}(s)$ (top) and $\hat{W}_{0}/\hat{W}_{0}^{\rm CMASS}$ (bottom) in the range of $s\in[15,57]~{}h^{-1}{\rm Mpc}$, which are relevant for $\chi^{2}$ calculations (Eq. 7). The error bars represent $2\sigma$ errors estimated by the PATCHY mocks. COLA and GADGET mocks exhibit good agreement under $\Omega_{m}=0.2951$, matching CMASS data within about $2\sigma$ level. For all four cases, there is one point slightly exceeding $2\sigma$ in both COLA and GADGET, which can be explained by the cosmic variance.
To further validate the consistency of the COLA mock, we conducted a comparison of the normalized 2PCF and MCFs derived from the CMASS-NGC data, GADGET, and PATCHY. This comparison focuses on larger clustering scales within the range of $s\in[15,57]~{}h^{-1}{\rm Mpc}$, which are relevant for the $\chi^{2}$ calculations [refer to Eq. (7)]. The results are shown in Fig. 12 to illustrate the normalized MCFs, $s^{2}\hat{W}_{0}(s)$ (top) and $\hat{W}_{0}/\hat{W}_{0}^{\rm CMASS}$ (bottom). The error bars represent the $2\sigma$ errors, estimated using the PATCHY mocks.
It is evident that the normalized statistics for the COLA and GADGET mocks exhibit good agreement under the same cosmological parameter $\Omega_{m}=0.2951$. Both COLA and GADGET simulations closely match the CMASS data within approximately $2\sigma$ level. These statistics from PATCHY are also in agreement with those from the CMASS data at the $2\sigma$ level. Both the COLA and GADGET simulations show one point beyond $2\sigma$ in all four cases, essentially due to the cosmic variance. Thus, we observe a good consistency between COLA and the high-accuracy GADGET simulations, indicating that COLA adequately ensures the reliability of our analysis.
For the stage-III surveys such as SDSS, the accuracy of COLA is sufficient given the associated statistical error of the data sample. However, for the ongoing stage-IV surveys, the statistical error of the sample is greatly reduced. In this case, it is necessary to improve the work of Ding et al. (2024) to ensure that the COLA achieves a higher degree of accuracy. Alternatively, one can directly utilize large-scale, high-accuracy cosmological N-body simulations that match the stage-IV surveys such as AbacusSummit Maksimova et al. (2021) and Uchuu Ishiyama et al. (2021).